# What Do Temporal Graph Learning Models Learn?

### v2

@@ 1 Introduction
Learning on temporal graphs has become an increasingly popular research topic, exemplified by the emergence of a multitude of benchmarks [[, e.g., ]]huang_temporal_2023, huang_benchtemp_2024, gastinger_tgb_2024, on which state-of-the-art graph learning models often appear to achieve very strong results. At the same time, the benchmark performances on link prediction tasks have faced increased scrutiny: from flaws in test sets and evaluation metrics leading to unrealistic results [undefq, undefo, undefc] to heuristics, such as predicting edges involving recently active and globally popular nodes [undefd], performing on par with state-of-the-art models. Further, even if specific models perform well on given benchmarks, it is not clear which factors contribute to this or, more specifically, which graph characteristics models pick up on to form their predictions.
@@ 2.1 Preliminaries
Temporal Graphs. In literature, temporal graphs are typically modeled either as continuous-time or discrete-time graphs [undefab, undefk].
@@
Datasets can be transformed between the two graph settings, although transformation from continuous to discrete-time typically causes some information loss when binning highly granular timestamps to discrete timesteps [undefk]. For the scope of this work, we focus on models designed for continuous-time dynamic graphs, as these are now more commonly used within the research community. However, we will often consider datasets with discretized timesteps due to our specific experimental designs. Within the scope of this work, unless otherwise specified, we generally ignore the use of edge and node features to simplify notation.
Representation Learning for Temporal Link Prediction. Most state-of-the-art graph learning models for dynamic graphs are based on different kinds of neural networks, from recurrent neural networks [undefl], to graph neural networks [undefs] and graph transformers [undefaa]. During training, continuous-time models receive the sequence of edges up to a time as input and learn a time-aware representation for each node . For dynamic link prediction, the input edges are directly used as positive examples, and optimized such that the existence of an edge at test time can be predicted from the corresponding pair of representations . There is some technical complexity in this optimization, as most models also require negative, i.e., non-existent edges, in training. Using all negative edges is typically infeasible since the size of the training data grows quadratically with the number of nodes. Further, empirical social networks are typically very sparse, such that models would be prone to predicting all edges as negative. Instead, a set of negative training edges needs to be sampled [undefy]. This is further complicated in the temporal context by also needing to consider issues such as the timestamps of negative samples, whether or not to specifically sample edges which have been present previously and whether to include nodes which will only become active at a future timestamp. Similar issues also carry over to the model evaluation, where negative samples are incorporated in any test set. In the literature, it is often the case that, for each positive edge, a negative edge at the same timestamp is sampled, but there are more complex strategies such as explicitly sampling from previously positive edges [undefaa].
@@ 2.2 Related Work
Dynamic Graph Learning Benchmarks. While popular general graph learning benchmarks, such as Open Graph Benchmark [undefh], also include link prediction tasks with time-based splits, benchmarks specifically targeting temporal graphs have emerged in recent years. Most notably, Temporal Graph Benchmark (TGB) [undefj] and BenchTemp [undefi] evaluate temporal graph learning models on node, link and graph classification tasks, showing that although strong results can be achieved on most datasets, there is no single solution across distinct datasets. TGB was later extended to heterogeneous and knowledge graphs by [undefe]. [undefz] identified limitations of temporal graph learning models in capturing sequential patterns in data, and established a benchmark focused on such data, while [undeff] built separate benchmarks for spatio-temporal, discrete-time and continuous-time graph datasets. The distinction between discrete and continuous-time models is bridged by [undefk], who present a unified framework and find that, despite some information loss, discrete-time models yield competitive results compared to continuous-time models, while being orders of magnitude faster at inference.
Limitations of Temporal Link Prediction Models. Several works have identified limitations in link prediction evaluation and the predictive capabilities of temporal graph learning models. [undefq] found that strong performances in dynamic link prediction is often explained by task simplicity, as evidenced by their EdgeBank baseline. EdgeBank predicts a positive edge at test time if and only if it was observed in training, yielding similar accuracy to state-of-the-art methods on several datasets. Thus, EdgeBank was also included in benchmarks such as TGB [undefj] and BenchTemp [undefi]. [undefn] further demonstrated that batch-based evaluation protocols can skew reported performance. [undefr] revealed limitations of temporal graph learning models in learning temporal patterns, as perturbations such as repeating positive edges with slightly altered timestamps or shuffling timestamps among training edges often have minimal impact on performance. [undefd] showed that simple heuristics favoring links to popular or recently active nodes can outperform state-of-the-art models on several datasets from TGB and BenchTemp. Further, they identified flaws in evaluation practices: commonly used rank-based metrics computed on sampled edge sets can yield results inconsistent with rankings on the full edge sets. In a broader context, [undefa] warned that poor benchmarking could lead to graph learning as a field losing relevance.
@@ 3 Experimental Framework
In Section 4, for each studied characteristic we systematically introduce the approach behind our evaluation, and follow it immediately with the findings. Within this section, we discuss experimental decisions that impact multiple characteristic experiments.
Datasets. In our experiments, we selected datasets specific to each characteristic, using both empirical and synthetic data. As empirical datasets, we chose the Bitcoin-Alpha [undefm], Enron [undeft], UCI [undefp] and Wikipedia [undefl] datasets, due to both their popularity in related work and computationally reasonable size in light of the extensiveness of our experiments. For these datasets, we follow [undefk] for discretizing the timestamps, going from UNIX timestamps to monthly, weekly or hourly granularity. The impact of this discretization is explored as the first characteristic under temporal granularity in our experiments (see Section 4.1). More details on the datasets are provided in Appendix A.1. For the temporal granularity, direction and density characteristics, the empirical datasets were split into into training, validation, and test data based on these discretized timesteps, with the first 70% of times being used as training data, and the remaining 30% evenly split into validation and test data. For the other characteristics, we designed synthetic graphs with direct correspondence to the studied characteristic and present these in the relevant results section.
Models and Training Setup. In our framework, we consider DyGFormer [undefaa], GraphMixer [undefb], DyRep [undefu], JODIE [undefl], TGN [undefs], TCL [undefv], TGAT [undefx] and CAWN [undefw], using the implementations from DyGLib [undefaa]. As hyperparameters, we used fixed values which gave effective performance on the empirical datasets. The values are set out in Appendix A.2, together with details of limited hyperparameter tuning in Appendix C. Unless noted otherwise, models were trained with accuracy on validation data used as early stopping criterion. Where validation data was not logically related to the training data, the training loss was used instead. We trained models ten times on each dataset, using different training seeds, or five times for each of five generation seeds for synthetic datasets. When not using the benchmark evaluation, testing is based on predictions for all possible edges at test time.
@@ 4.1.1 Temporal Granularity
Discrete: the discrete-time variants of the datasets, following [undefk].
@@ 4.2.2 Periodicity
Findings. Results for periodicity in terms of confidence scores on the UCI dataset are depicted in Figure 4. These are largely in line with the results for other datasets, period lengths, and the corresponding accuracy scores in Appendix B. We observe some differences compared to our experiments on persistence. Most notably, CAWN and DyGFormer do not appear to distinguish between odd and even timesteps, and simply give high confidence to every edge that occurs at all (✗). This behavior is very similar to how the EdgeBank baseline [undefq] would perform. Similarly, DyRep does not appear to distinguish between odd and even timesteps (✗). By contrast, GraphMixer and TGAT appear to distinguish odd and even timesteps very well (✓). Finally, JODIE, TGAT and TCL appear to consistently separate edges from different timesteps, but tend to be too confident in predictions on edges from the wrong timestep ().
@@ 4.2.3 Recency
For all deep learning models, it is often anticipated that models are more influenced by the latest data in training. This is even more of a consideration when the data itself contains a temporal element, as with dynamic graphs. Recent work has also shown that heuristics which emphasize predicting edges on more recently active nodes can perform very strongly on graph datasets [undefd]. Thus, we look to evaluate the extent to which model predictions are biased toward more recently seen data.
@@
Findings. In Figure 5, we observe that, across all models and datasets, average probability scores of seen edges do not appear to vary based on the time they were last observed. Instead, they stay largely constant, with the exception of TGAT which appears to have some bias toward earlier edges on the Wikipedia-based graphs. This indicates that, in general, models do not place any predictive weight into the recency of edges (✗). Since edges were sampled completely at random, there is no additional information for models based on which predictions could be made and models consequently display high uncertainty in their probability scores. Given that heuristics based on predicting edges to recently active nodes often perform on par with state-of-the-art models [undefd], we argue that this yields potential room for improvement of this state-of-the-art. Introducing parameters that control whether a model places more weight on recent observations could improve performance on many datasets.
@@ 4.3.1 Homophily
Approach. To examine whether the models learn homophily, we use homophilic and heterophilic stochastic block models (SBMs) [undefg] with 1000 nodes, split into two groups of 500. For homophilic SBMs, we set the likelihood of inter-group links to and intra-group links to : reversed for heterophilic SBMs. Based on these parameters, we created dynamic networks of timesteps by resampling edges times. Intuitively, if a model learns homophily, it should mimic the edge formation of the homophilic SBM by placing higher likelihood on predicting intra-group rather than inter-group edges, with intra-group links for both groups being equally likely—showing opposite behavior for heterophilic SBMs. We test this assumption by aggregating predicted probabilities of edges within and between groups. During training, the node groups are explicitly provided as one-hot-encoded node features.
@@ 4.3.2 Preferential Attachment
Approach. We create graphs based on the Barabási-Albert (BA) model [undef] with 1000 nodes and 2000 edges at each timestep, for 100 training and 21 validation timesteps. The initial timestep is generated with the standard BA model to introduce variation in node degree. Subsequent timesteps add edges by sampling two nodes proportional to their number of previous edges—see Appendix A.4 for a formalization. Evaluation is done by examining the relationship between each node’s degree and the average probabilities predicted for its potential edges which were never positive in training. If preferential attachment is learned, then the edges connecting to nodes of high degree should receive higher average probabilities.
@@ 5 Discussion and Conclusions
Contributions. We propose a novel framework that assesses how well graph learning models capture intuitive characteristics of temporal networks. Based on this framework, we identify limitations of existing models: such as a limited distinction between the directions of edges, failure to detect periodic patterns, or a lack of emphasis on more recently observed graph dynamics. We do not argue that, in general, temporal graph learning models must capture all of these characteristics, as there may be application scenarios in which individual characteristics are not desirable or necessary. For instance, it may sometimes be counterproductive to place predictive emphasis on more recently seen edges. However, understanding a model’s predictive capabilities is crucial for selecting the right one in practice. In that sense, our work can be seen as both a benchmark for a more interpretability-driven evaluation of temporal graph learning models and, complementing existing performance-oriented benchmarks [undefj, undefi, undefe, undefz], as an aid for researchers and practitioners when choosing a model for their task at hand.
@@ A.1 Details of Empirical Datasets
Enron [undeft] is a network modeling email communication within the Enron corporation between 2000 and 2002. Nodes correspond to individuals that worked at Enron, and an edge indicates that an email was sent from individual to individual at time .
@@
Bitcoin-Alpha [undefm] is a whom-trust-whom network of bitcoin users that traded on the platform http://www.btc-aplha.com. An edge indicates that user assigned a rating to user at time . Each edge further contains the rating score which ranges from -10 to 10.
@@
UCI [undefp] represents an online community of students at the University of California, Irvine. Within this community, students could search each other’s profiles and message each other. The network is then modeled based on messages sent in the period from April to October 2004. An edge indicates that user sent a message to user at time .
@@
Wikipedia [undefl] is a heterogeneous network that models one month of edits on Wikipedia pages. An edge indicates that editor edited page at time .
@@ A.2 Choice of Hyperparameters
For the datasets and models we used, [undefaa] found recent sampling of neighbors to always be the most effective. There was more variation in the other parameters depending on the dataset and model combination. The values fixed here give very similar Average Precision scores when following the DyGLib evaluation procedure compared to the optimal dataset specific hyperparameters specified in the paper (see Table 5).

### v3

@@ 1 Introduction
Learning on temporal graphs has become an increasingly popular research topic, exemplified by the emergence of a multitude of benchmarks [[, e.g., ]]huang_temporal_2023, huang_benchtemp_2024, gastinger_tgb_2024, on which state-of-the-art graph learning models often appear to achieve very strong results. At the same time, the benchmark performances on link prediction tasks have faced increased scrutiny: from flaws in test sets and evaluation metrics leading to unrealistic results [18, 16, 4] to heuristics, such as predicting edges involving recently active and globally popular nodes [5], performing on par with state-of-the-art models. Further, even if specific models perform well on given benchmarks, it is not clear which factors contribute to this or, more specifically, which graph characteristics models pick up on to form their predictions.
@@ 2.1 Preliminaries
Temporal Graphs. In literature, temporal graphs are typically modeled either as continuous-time or discrete-time graphs [29, 12].
@@
Datasets can be transformed between the two graph settings, although transformation from continuous to discrete-time typically causes some information loss when binning highly granular timestamps to discrete timesteps [12]. For the scope of this work, we focus on models designed for continuous-time dynamic graphs, as these are now more commonly used within the research community. However, we will often consider datasets with discretized timesteps due to our specific experimental designs. Within the scope of this work, unless otherwise specified, we generally ignore the use of edge and node features to simplify notation.
Representation Learning for Temporal Link Prediction. Most state-of-the-art graph learning models for dynamic graphs are based on different kinds of neural networks, from recurrent neural networks [14], to graph neural networks [20] and graph transformers [28]. During training, continuous-time models receive the sequence of edges up to a time as input and learn a time-aware representation for each node . For dynamic link prediction, the input edges are directly used as positive examples, and optimized such that the existence of an edge at test time can be predicted from the corresponding pair of representations . There is some technical complexity in this optimization, as most models also require negative, i.e., non-existent edges, in training. Using all negative edges is typically infeasible since the size of the training data grows quadratically with the number of nodes. Further, empirical social networks are typically very sparse, such that models would be prone to predicting all edges as negative. Instead, a set of negative training edges needs to be sampled [26]. This is further complicated in the temporal context by also needing to consider issues such as the timestamps of negative samples, whether or not to specifically sample edges which have been present previously and whether to include nodes which will only become active at a future timestamp. Similar issues also carry over to the model evaluation, where negative samples are incorporated in any test set. In the literature, it is often the case that, for each positive edge, a negative edge at the same timestamp is sampled, but there are more complex strategies such as explicitly sampling from previously positive edges [28].
@@ 2.2 Related Work
Dynamic Graph Learning Benchmarks. While popular general graph learning benchmarks, such as Open Graph Benchmark [9], also include link prediction tasks with time-based splits, benchmarks specifically targeting temporal graphs have emerged in recent years. Most notably, Temporal Graph Benchmark (TGB) [11] and BenchTemp [10] evaluate temporal graph learning models on node, link and graph classification tasks, showing that although strong results can be achieved on most datasets, there is no single solution across distinct datasets. TGB was later extended to heterogeneous and knowledge graphs by [6]. [27] identified limitations of temporal graph learning models in capturing sequential patterns in data, and established a benchmark focused on such data, while [7] built separate benchmarks for spatio-temporal, discrete-time and continuous-time graph datasets. The distinction between discrete and continuous-time models is bridged by [12], who present a unified framework and find that, despite some information loss, discrete-time models yield competitive results compared to continuous-time models, while being orders of magnitude faster at inference.
Limitations of Temporal Link Prediction Models. Several works have identified limitations in link prediction evaluation and the predictive capabilities of temporal graph learning models. [18] found that strong performances in dynamic link prediction is often explained by task simplicity, as evidenced by their EdgeBank baseline. EdgeBank predicts a positive edge at test time if and only if it was observed in training, yielding similar accuracy to state-of-the-art methods on several datasets. Thus, EdgeBank was also included in benchmarks such as TGB [11] and BenchTemp [10]. [15] further demonstrated that batch-based evaluation protocols can skew reported performance. [19] revealed limitations of temporal graph learning models in learning temporal patterns, as perturbations such as repeating positive edges with slightly altered timestamps or shuffling timestamps among training edges often have minimal impact on performance. [5] showed that simple heuristics favoring links to popular or recently active nodes can outperform state-of-the-art models on several datasets from TGB and BenchTemp. Further, they identified flaws in evaluation practices: commonly used rank-based metrics computed on sampled edge sets can yield results inconsistent with rankings on the full edge sets. In a broader context, [2] warned that poor benchmarking could lead to graph learning as a field losing relevance.
@@ 3 Experimental Framework
In Section 4, for each studied characteristic we systematically introduce the approach behind our evaluation, and follow it immediately with the findings. Within this section, we discuss experimental decisions that impact multiple characteristic experiments. Further, we provide the code used for our experiments on https://github.com/dess-mannheim/tgl.
Datasets. In our experiments, we selected datasets specific to each characteristic, using both empirical and synthetic data. As empirical datasets, we chose the Bitcoin-Alpha [13], Enron [21], UCI [17] and Wikipedia [14] datasets, due to both their popularity in related work and computationally reasonable size in light of the extensiveness of our experiments. For these datasets, we follow [12] for discretizing the timestamps, going from UNIX timestamps to monthly, weekly or hourly granularity. The impact of this discretization is explored as the first characteristic under temporal granularity in our experiments (see Section 4.1). More details on the datasets are provided in Appendix A.1. For the temporal granularity, direction and density characteristics, the empirical datasets were split into into training, validation, and test data based on these discretized timesteps, with the first 70% of times being used as training data, and the remaining 30% evenly split into validation and test data. For the other characteristics, we designed synthetic graphs with direct correspondence to the studied characteristic and present these in the relevant results section.
Models and Training Setup. In our framework, we consider DyGFormer [28], GraphMixer [3], DyRep [22], JODIE [14], TGN [20], TCL [23], TGAT [25] and CAWN [24], using the implementations from DyGLib [28]. As hyperparameters, we used fixed values which gave effective performance on the empirical datasets. The values are set out in Appendix A.2, together with details of limited hyperparameter tuning in Appendix C. Unless noted otherwise, models were trained with accuracy on validation data used as early stopping criterion. Where validation data was not logically related to the training data, the training loss was used instead. We trained models ten times on each dataset, using different training seeds, or five times for each of five generation seeds for synthetic datasets. When not using the benchmark evaluation, testing is based on predictions for all possible edges at test time.
@@ 4.1.1 Temporal Granularity
Discrete: the discrete-time variants of the datasets, following [12].
@@ 4.2.2 Periodicity
Findings. Results for periodicity in terms of confidence scores on the UCI dataset are depicted in Figure 4. These are largely in line with the results for other datasets, period lengths, and the corresponding accuracy scores in Appendix B. We observe some differences compared to our experiments on persistence. Most notably, CAWN and DyGFormer do not appear to distinguish between odd and even timesteps, and simply give high confidence to every edge that occurs at all (✗). This behavior is very similar to how the EdgeBank baseline [18] would perform. Similarly, DyRep does not appear to distinguish between odd and even timesteps (✗). By contrast, GraphMixer and TGAT appear to distinguish odd and even timesteps very well (✓). Finally, JODIE, TGAT and TCL appear to consistently separate edges from different timesteps, but tend to be too confident in predictions on edges from the wrong timestep ().
@@ 4.2.3 Recency
For all deep learning models, it is often anticipated that models are more influenced by the latest data in training. This is even more of a consideration when the data itself contains a temporal element, as with dynamic graphs. Recent work has also shown that heuristics which emphasize predicting edges on more recently active nodes can perform very strongly on graph datasets [5]. Thus, we look to evaluate the extent to which model predictions are biased toward more recently seen data.
@@
Findings. In Figure 5, we observe that, across all models and datasets, average probability scores of seen edges do not appear to vary based on the time they were last observed. Instead, they stay largely constant, with the exception of TGAT which appears to have some bias toward earlier edges on the Wikipedia-based graphs. This indicates that, in general, models do not place any predictive weight into the recency of edges (✗). Since edges were sampled completely at random, there is no additional information for models based on which predictions could be made and models consequently display high uncertainty in their probability scores. Given that heuristics based on predicting edges to recently active nodes often perform on par with state-of-the-art models [5], we argue that this yields potential room for improvement of this state-of-the-art. Introducing parameters that control whether a model places more weight on recent observations could improve performance on many datasets.
@@ 4.3.1 Homophily
Approach. To examine whether the models learn homophily, we use homophilic and heterophilic stochastic block models (SBMs) [8] with 1000 nodes, split into two groups of 500. For homophilic SBMs, we set the likelihood of inter-group links to and intra-group links to : reversed for heterophilic SBMs. Based on these parameters, we created dynamic networks of timesteps by resampling edges times. Intuitively, if a model learns homophily, it should mimic the edge formation of the homophilic SBM by placing higher likelihood on predicting intra-group rather than inter-group edges, with intra-group links for both groups being equally likely—showing opposite behavior for heterophilic SBMs. We test this assumption by aggregating predicted probabilities of edges within and between groups. During training, the node groups are explicitly provided as one-hot-encoded node features.
@@ 4.3.2 Preferential Attachment
Approach. We create graphs based on the Barabási-Albert (BA) model [1] with 1000 nodes and 2000 edges at each timestep, for 100 training and 21 validation timesteps. The initial timestep is generated with the standard BA model to introduce variation in node degree. Subsequent timesteps add edges by sampling two nodes proportional to their number of previous edges—see Appendix A.4 for a formalization. Evaluation is done by examining the relationship between each node’s degree and the average probabilities predicted for its potential edges which were never positive in training. If preferential attachment is learned, then the edges connecting to nodes of high degree should receive higher average probabilities.
@@ 5 Discussion and Conclusions
Contributions. We propose a novel framework that assesses how well graph learning models capture intuitive characteristics of temporal networks. Based on this framework, we identify limitations of existing models: such as a limited distinction between the directions of edges, failure to detect periodic patterns, or a lack of emphasis on more recently observed graph dynamics. We do not argue that, in general, temporal graph learning models must capture all of these characteristics, as there may be application scenarios in which individual characteristics are not desirable or necessary. For instance, it may sometimes be counterproductive to place predictive emphasis on more recently seen edges. However, understanding a model’s predictive capabilities is crucial for selecting the right one in practice. In that sense, our work can be seen as both a benchmark for a more interpretability-driven evaluation of temporal graph learning models and, complementing existing performance-oriented benchmarks [11, 10, 6, 27], as an aid for researchers and practitioners when choosing a model for their task at hand.
@@
Acknowledgements
This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Grant No. 453349072 and Grant No. 530081187. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and DFG through grant INST 35/1597-1 FUGG.
@@ A.1 Details of Empirical Datasets
Enron [21] is a network modeling email communication within the Enron corporation between 2000 and 2002. Nodes correspond to individuals that worked at Enron, and an edge indicates that an email was sent from individual to individual at time .
@@
Bitcoin-Alpha [13] is a whom-trust-whom network of bitcoin users that traded on the platform http://www.btc-aplha.com. An edge indicates that user assigned a rating to user at time . Each edge further contains the rating score which ranges from -10 to 10.
@@
UCI [17] represents an online community of students at the University of California, Irvine. Within this community, students could search each other’s profiles and message each other. The network is then modeled based on messages sent in the period from April to October 2004. An edge indicates that user sent a message to user at time .
@@
Wikipedia [14] is a heterogeneous network that models one month of edits on Wikipedia pages. An edge indicates that editor edited page at time .
@@ A.2 Choice of Hyperparameters
For the datasets and models we used, [28] found recent sampling of neighbors to always be the most effective. There was more variation in the other parameters depending on the dataset and model combination. The values fixed here give very similar Average Precision scores when following the DyGLib evaluation procedure compared to the optimal dataset specific hyperparameters specified in the paper (see Table 6).
@@ A.4 Preferential Attachment Model
| Bitcoin-Alpha | Enron | UCI | Wikipedia | |||||||||||||
| Cont. | Disc. | Flat | Cont. | Disc. | DF | Flat | FF | Cont. | Disc. | DF | Flat | FF | Cont. | Disc. | Flat | |
| CAWN | 0.780 | 0.984 | 0.728 | 0.959 | 0.941 | 0.903 | 0.518 | 0.630 | 0.968 | 0.953 | 0.971 | 0.684 | 0.671 | 0.989 | 0.983 | 0.571 |
| DyGF. | 0.862 | 0.989 | 0.724 | 0.952 | 0.939 | 0.922 | 0.751 | 0.639 | 0.964 | 0.960 | 0.972 | 0.532 | 0.463 | 0.988 | 0.980 | 0.495 |
| DyRep | 0.893 | 0.905 | 0.941 | 0.891 | 0.732 | 0.830 | 0.576 | 0.810 | 0.925 | 0.889 | 0.935 | 0.834 | 0.916 | 0.965 | 0.943 | 0.829 |
| Mixer | 0.997 | 0.993 | 0.439 | 0.951 | 0.871 | 0.823 | 0.451 | 0.459 | 0.983 | 0.917 | 0.944 | 0.477 | 0.518 | 0.975 | 0.945 | 0.435 |
| JODIE | 0.999 | 0.999 | 0.961 | 0.936 | 0.860 | 0.911 | 0.701 | 0.853 | 0.958 | 0.946 | 0.964 | 0.845 | 0.941 | 0.969 | 0.955 | 0.829 |
| TCL | 0.880 | 0.991 | 0.668 | 0.842 | 0.870 | 0.832 | 0.580 | 0.554 | 0.956 | 0.949 | 0.969 | 0.624 | 0.677 | 0.971 | 0.952 | 0.549 |
| TGAT | 0.838 | 1.000 | 0.582 | 0.756 | 0.875 | 0.844 | 0.496 | 0.488 | 0.869 | 0.926 | 0.925 | 0.463 | 0.491 | 0.968 | 0.960 | 0.506 |
| TGN | 0.995 | 0.999 | 0.955 | 0.908 | 0.889 | 0.912 | 0.598 | 0.904 | 0.982 | 0.948 | 0.971 | 0.794 | 0.915 | 0.985 | 0.971 | 0.848 |